A ball thrown vertically upwards with aspeed of 19.6m/s from thetop of tower returns to the earth in 6 s.Find the height of the tower
 

Grade:11

2 Answers

Himanshu Rathour
5 years ago
Time taken to reach highest point
t=u/g
t=19.6/9.8
t=2s
Maximum height reached h=u²/2g
=19.6
Total time of flight =6s
Time of ascent=2s
Thus 
Time of descent =4s
Height covered in reaching the earth
H=1/2g(16)
=78.4
Height of tower=H-h
=58.8m
Khimraj
5 years ago
I am taking downward direction as positive and upward as negative.
h = – ut +½*gt2
h = – 19.6*6 +  ½*9.8*62
h = 19.6*6(1.5 – 1) = 58.8 m
Hope it clears.